Solve (a/x)-(y/b)=3 for x.
Answer by sachi(548)

(a/x)-(y/b)=3
or (a/x)=(y/b)+3=(y+3b)/3
or x/a=3/(y+3b)
or x=3a/(y+3b)
ans